### v1

# H-CMRH: A Novel Inner Product Free Hybrid Krylov Method for Large-Scale Inverse Problems

Abstract
This study investigates the iterative regularization properties of two Krylov methods for solving large-scale ill-posed problems: the changing minimal residual Hessenberg method (CMRH) and a novel hybrid variant called the hybrid changing minimal residual Hessenberg method (H-CMRH). Both methods share the advantages of avoiding inner products, making them efficient and highly parallelizable, and particularly suited for implementations that exploit randomization and mixed precision arithmetic. Theoretical results and extensive numerical experiments suggest that H-CMRH exhibits comparable performance to the established hybrid GMRES method in terms of stabilizing semiconvergence, but H-CMRH has does not require any inner products, and requires less work and storage per iteration.
keywords:
inverse problems, Tikhonov regularization,
{MSCcodes}
AMS Subject Classifications: 65F22, 65F10, 49N45, 65K99
1 Introduction
Linear large-scale discrete inverse problems can be written in the form
$Ax_{{true}}+e=b\,,$ (1)
where $A\in\mathbb{R}^{n\times n}$ models the forward problem, $x_{{true}}\in\mathbb{R}^{n}$ is the unknown solution we want to approximate, $b\in\mathbb{R}^{n}$ is the vector of measurements, and $e\in\mathbb{R}^{n}$ represents errors in the measured data, typically assumed to come from independent and identically distributed (iid) zero mean normal random variables (so-called Gaussian white noise). Inverse problems of the form (1) arise in a variety of applications including medical and geophysical imaging, and image deblurring [10, 19, 3, 20]. Solving (1) is cumbersome due to its ill-posed nature, meaning that the singular values of the system matrix $A$ decay and cluster at zero without a gap to indicate numerical rank. This property, namely $A$ being moderately to severely ill-conditioned, causes the solution to be sensitive to perturbations in the measurement vector $b$; that is, small changes in $b$ can produce large changes in the estimate of the true solution $x_{{true}}$.
Regularization is typically used to alleviate this sensitivity [10] and enables the production of meaningful approximations of the true solution. For example, a very relevant approach when dealing with large-scale problems is to employ iterative regularization, which consists of applying an iterative solver to the least squares problem
$\min_{x\in\mathbb{R}^{n}}\|b-Ax\|_{2}^{2}\,,$ (2)
where regularization is achieved through early termination of the iterations [2]. The stopping iteration acts a regularization parameter: if the iterative solver is terminated too soon, the solution is overly smooth, whereas the solution is highly oscillatory if the iterative solver is terminated too late.
Alternatively, one can use a variational approach such as Tikhonov regularization, which consists of solving an optimization problem of the form
$\min_{x\in\mathbb{R}^{n}}\|b-Ax\|_{2}^{2}+\lambda^{2}\|Wx\|_{2}^{2}\,,$ (3)
where $\lambda$ is the regularization parameter and $\|Wx\|_{2}^{2}$ with $W\in\mathbb{R}^{n\times n}$ is the regularization term [2]. Similarly to the role of the stopping iteration in iterative regularization, it is important to choose a good value for $\lambda$; if it is chosen too large, the regularized solution is overly smooth, and a choice of $\lambda$ that is too small produces a highly oscillatory solution. Moreover, multiple linear solves might be needed to refine the choice of $\lambda$, drastically increasing the computational cost. Thus, when dealing with large scale problems, it may be necessary to solve (3) with an iterative method.
Hybrid regularization is a particular combination of variational and iterative regularization. This approach consists of iteratively projecting (2) onto a small subspace of increasing dimension and applying variational (e.g., Tikhonov) regularization to the small projected problem. In this framework, we only require matrix-vector products with $A$, which allows us to avoid explicitly constructing or storing $A$. It also creates a natural environment, namely the small projected problems, for estimating a good regularization parameter [2].
In this study, we analyze the iterative regularization properties of the changing minimal residual Hessenberg method (CMRH) and introduce a new hybrid projection algorithm known as the hybrid changing minimal residual Hessenberg method (H-CMRH). Both CMRH and H-CMRH are Krylov methods that avoid inner products, making them very appealing for modern computations since they are very efficient and highly paralellizable. In particular, these characteristics make them very suitable for randomized methods and for mixed precision arithmetic.
[…]
2 The Changing Minimal Residual Hessenberg Method
Both CMRH in [15] and GMRES [14] iteratively approximate the solution of (2) on a Krylov subspace of increasing dimension:
${\cal{K}}_{k}=span\{r_{0},Ar_{0},A^{2}r_{0},.....,A^{k-1}r_{0}\}$
where $r_{0}=b-Ax_{0}$ and $x_{0}$ is an initial guess for the solution.
Within each of these two iterative solvers, different basis vectors of ${\cal{K}}_{k}$ are formed. On one hand, CMRH utilizes the Hessenberg process which produces linearly independent basis vectors. On the other hand, for GMRES, the Arnoldi algorithm forms orthonormal basis vectors of ${\cal{K}}_{k}$. Despite the lack of orthonormal vectors, CMRH is still advantageous. For dense problems, the matrix-vector product with $A$ can be calculated without accessing the first $k$ columns of $A$. This is a result of the basis vectors containing zeros in the first $k$--$1$ entries, and thus giving possible savings in the cost of the matrix-vector product [17]. More generally, CMRH does not require inner products, which is important for computer efficiency, parallel computing and possibly reducing errors caused by inner products in low precision.
2.1 The Hessenberg Process and CMRH
Consider the following Krylov matrix:
$V_{k}=[r_{0},Ar_{0},A^{2}r_{0},.........,A^{k-1}r_{0}]\in\mathbb{R}^{n\times k},$ (4)
where $V_{k+1}=[r_{0},\,AV_{k}]$ and clearly range$(V_{k})={\cal{K}}_{k}$. Note that $V_{k}$ becomes very ill-conditioned even for small numbers of $k$, so this matrix is not constructed explicitly, and is only considered theoretically to motivate the Hessenberg process, and later on its relationship with the Arnoldi process.
The Hessenberg process can be derived by considering the LU factorization of
$V_{k}=L_{k}U_{k},$ (5)
where $L_{k}\in\mathbb{R}^{n\times k}$ is a unit lower triangular matrix and $U_{k}$ is an upper triangular matrix. However, the algorithm does not explicitly compute this factorization but instead recursively computes the columns of $L_{k}$. Following [17] we can write the following relation:
$V_{k+1}\begin{bmatrix}0_{1\times k}\\ I_{k}\end{bmatrix}=L_{k+1}U_{k+1}\begin{bmatrix}0_{1\times k}\\ I_{k}\end{bmatrix}=AV_{k}=AL_{k}U_{k},$ (6)
where $0_{1\times k}$ is a row vector of zeros of dimensions $1$$\times$$k$. One can then define an upper Hessenberg matrix of the form:
$H_{{k+1},k}=U_{k+1}\begin{bmatrix}0_{1\times k}\\ I_{k}\end{bmatrix}U_{k}^{-1}\in\mathbb{R}^{(k+1)\times k},$ (7)
where $k<n$. Note that the columns of $L_{k}$ are linearly independent vectors and form a basis of the Krylov subspace ${\cal{K}}_{k}$. Moreover, combining (6) and (7) leads to the following Hessenberg relation:
$AL_{k}=L_{k+1}H_{{k+1},k}.$ (8)
An implementation of the Hessenberg process [16] can be found in Algorithm 1. The algorithm requires providing a nonzero vector $r_{0}$; we discuss below the specific choice for this vector in the CMRH method.
Algorithm 1 Hessenberg Process
1:$\beta=e_{1}^{T}r_{0}$ (first entry of given, nonzero vector $r_{0}$); $l_{1}=r_{0}/\beta$
2:for $k=1,....,m$ do
3:     $u=Al_{k}$
4:     for $j=1,....,k$ do
5:         $h_{j,k}=u(j)$; $u=u-h_{j,k}l_{j}$
6:     end for
7:     $h_{k+1,k}=u(k+1)$; $l_{k+1}=u/h_{k+1,k}$
8:end for
Examining this implementation, we find that if $\beta=0$ or $h_{k+1,k}=0$, then the Hessenberg process will breakdown. To circumvent this breakdown, a pivoting approach is introduced in [16]. The Hessenberg process with pivoting is described in Algorithm 2.
Algorithm 2 Hessenberg Process with Pivoting
[…]
2:Determine $i_{0}$ such that $|r_{0}(i_{0})|=\|r_{0}\|_{\infty}$
3:$\beta=r_{0}(i_{0})$ ; $l_{1}=r_{0}/\beta$; $p_{1}\Leftrightarrow p_{i_{0}}$
4:for $k=1,....,m$ do
5:     $u=Al_{k}$
6:     for $j=1,....,k$ do
7:         $h_{j,k}=u(p(j))$ ; $u=u-h_{j,k}l_{j}$
8:     end for
9:     if $k<n$ and $u\neq 0$ then
10:         Determine $i_{0}\in\{k+1,.......,n\}$ such that $|u(p(i_{0}))|=\|u(p(k+1:n))\|_{\infty}$
11:         $h_{k+1,k}=u(p(i_{0}))$; $l_{k+1}=u/h_{k+1,k}$ ; $p_{k+1}\Leftrightarrow p_{i_{0}}$
12:     else
13:         $h_{k+1,k}=0$; Stop.
14:     end if
15:end for
At each iteration $k$, CMRH finds an approximate solution of (2) by minimizing the following least squares problem:
$\min_{x\in{\cal{K}}_{k}}\|L^{\dagger}_{k+1}(b-Ax)\|,$ (9)
where $L^{\dagger}_{k+1}$ is the pseudoinverse of the lower triangular matrix $L_{k+1}$ (which is constructed with the Hessenberg process). We observe that CMRH is a projection method since, at each iteration, it considers a solution on the Krylov subspace ${\cal{K}}_{k}$. However, due to $L_{k+1}^{\dagger}$ not being orthogonal, the subproblems of the form (9) solved at each iteration are oblique projections of the original problem (2), and therefore CMRH is not exactly a minimal residual method. For that reason, CMRH holds some parallels with the quasi-minimal residual method (QMR), in that the imposed optimality conditions involve minimizing a semi-norm [5]. Considering $x=x_{0}+L_{k}y$, and $r_{0}=b-Ax_{0}$, the norm in (9) can be rewritten as:
$\|L^{\dagger}_{k+1}(b-A(x_{0}+L_{k}y))\|$ $=$ $\|L^{\dagger}_{k+1}(r_{0}-AL_{k}y)\|$
$=$ $\|L^{\dagger}_{k+1}(r_{0}-L_{k+1}H_{k+1,k}y)\|$
$=$ $\|\beta e_{1}-H_{k+1,k}y\|,$
where we have used the Hessenberg relation (7), and where $\beta$ is the first entry in $r_{0}$ (if using Algorithm 1) or the largest entry in $r_{0}$ (if using Algorithm 2). Thus, at each iteration, we solve a small subproblem on a space of increasing dimension
$y_{k}=\arg\min_{y\in\mathbb{R}^{k}}\|\beta e_{1}-H_{k+1}y\|$
and then project back using $x_{k}=x_{0}+L_{k}y$.
2.2 The Relationship Between CMRH and GMRES
Similar to CMRH, one can derive the GMRES method by considering the QR factorization of the Krylov matrix defined in (4):
$V_{k}=Q_{k}\tilde{R}_{k},$ (10)
where $Q_{k}\in\mathbb{R}^{n\times k}$ has orthonormal columns and $\tilde{R}_{k}$ is an upper triangular matrix. This algorithm does not explicitly compute the QR factorization but instead recursively computes the columns of $Q_{k}$. From [17] we consider the following relationship:
$V_{k+1}\begin{bmatrix}0_{1\times k}\\ I_{k}\end{bmatrix}=Q_{k+1}\tilde{R}_{k+1}\begin{bmatrix}0_{1\times k}\\ I_{k}\end{bmatrix}=AV_{k}=AQ_{k}\tilde{R}_{k},$ (11)
where $0_{1\times k}$ is a row vector of zeros of dimensions $1$$\times$$k$ and which produces an upper Hessenberg matrix of the form:
$H_{k+1,k}^{A}=\tilde{R}_{k+1}\begin{bmatrix}0_{1\times k}\\ I_{k}\end{bmatrix}\tilde{R}_{k}^{-1}\in\mathbb{R}^{(k+1)\times k},$ (12)
where $k<n$. Note that the columns of $Q_{k}$ form an orthonormal basis of ${\cal{K}}_{k}$. Moreover, combining (11) and (12) creates the Arnoldi relation:
$A_{k}Q_{k}=Q_{k+1}H_{k+1,k}^{A}.$ (13)
The derivation of (13) and (8) naturally provides a mapping between the Arnoldi and Hessenberg bases. Consider the LU and QR factorizations of the Krylov matrix $V_{k}$ in (5) and (10), respectively, then we can define
$V_{k}=L_{k}U_{k}=Q_{k}\tilde{R}_{k}.$ (14)
Now, let $R_{k}=\tilde{R}_{k}U_{k}^{-1}$: this is a $k\times k$ upper triangular matrix. Using (14), the lower triangular matrix $L_{k}$ can be written as a composition of an orthogonal matrix and an upper right triangular matrix
$L_{k}=Q_{k}\tilde{R}_{k}U_{k}^{-1}.$
This corresponds to a QR factorization of $L_{k}$,
$L_{k}=Q_{k}R_{k}.$ (15)
In addition to (15), we can rewrite the Arnoldi relation in (13) as:
$AL_{k}R^{-1}_{k}=L_{k+1}R^{-1}_{k+1}H_{k+1,k}^{A}.$ (16)
Comparing the Hessenberg (8) and Arnoldi (16) relations, and following [17], we provide the following proposition :
Proposition 2.1.
Let $H_{k+1,k}$ and $H^{A}_{k+1,k}$ be the Hessenberg matrices associated to the Hessenberg and Arnoldi processes, respectively, at iteration $k$, then
$H_{k+1,k}=R^{-1}_{k+1}H_{k+1,k}^{A}R_{k},$
or, equivalently, $H^{A}_{k+1,k}=R_{k+1}H_{k+1,k}R_{k}^{-1}.$
Proposition 2.1 is used to establish residual bounds for CMRH in the following sense: the residual norm associated to the approximated solution provided by CMRH at each iteration is close to the residual norm associated to the solution provided by GMRES if the condition number of $R_{k+1}$ does not grow too quickly. This can be observed in the following theorem, originally proved in [17]. Here, we restate it and provide additional details of the proof in order to set the stage for an analogous reasoning involving the hybrid CMRH algorithm (H-CMRH) in Section 4.
Theorem 2.2.
Let $r_{k}^{G}$ and $r_{k}^{C}$ be the GMRES and CMRH residuals at the kth iteration beginning with the same initial residual $r_{0}$, respectively. Then
$\|r_{k}^{G}\|\leq\|r_{k}^{C}\|\leq\kappa(R_{k+1})\|r_{k}^{G}\|$ (17)
where $\kappa(R_{k+1})=\|R_{k+1}\|\|R_{k+1}^{-1}\|$ is the condition number of $R_{k+1}$.
Proof 2.3.
First, we prove the left inequality in (17). Consider the residual as a function of the solution:
$r(x)=b-Ax.$
Then, the kth residual norm associated to the approximated solution produced by GMRES is:
$\|r_{k}^{G}\|=\|b-Ax_{k}^{G}\|=\min_{x\in{\cal{K}}_{k}}\|r(x)\|.$
Since $x_{k}^{G}$ and $x_{k}^{C}$ are in the Krylov subspace ${\cal{K}}_{k}$, then by definition:
$\min_{x\in{\cal{K}}_{k}}\|r(x)\|\leq\|r_{k}^{C}\|=\|r(x_{k}^{C})\|.$
Hence, $\|r_{k}^{G}\|\leq\|r_{k}^{C}\|$.
Now we prove the right inequality in (17). Since $r_{k}^{C}$ and $r_{k}^{G}$ are in the ${\cal{K}}_{k+1}$ subspace, we can write $r_{k}^{C}$ and $r_{k}^{G}$ as a linear combination of any basis of ${\cal{K}}_{k+1}$. Using the Hessenberg relation, the decomposition of the Krylov matrix is:
$V_{k+1}=L_{k+1}U_{k+1}.$
This implies range$(L_{k+1})$ = range$(V_{k+1})={\cal{K}}_{k+1}$. Therefore, and using (15), there exist $u_{k}^{C}$ and $w_{k}^{C}$ in $\mathbb{R}^{k+1}$ such that
$r_{k}^{C}=L_{k+1}u_{k}^{C}=Q_{k+1}R_{k+1}u_{k}^{C}=Q_{k+1}w_{k}^{C}$
with $R_{k+1}u_{k}^{C}=w_{k}^{C}$. Analogously, there exist $u_{k}^{G}$ and $w_{k}^{G}$ in $\mathbb{R}^{k+1}$ such that
$r_{k}^{G}=L_{k+1}u_{k}^{G}=Q_{k+1}R_{k+1}u_{k}^{G}=Q_{k+1}w_{k}^{G}$ (18)
with $R_{k+1}u_{k}^{G}=w_{k}^{G}$.
Consider the optimality conditions of CMRH. As stated above, $r_{k}^{C}=L_{k+1}u_{k}^{C}$. This implies that $L^{\dagger}_{k+1}r_{k}^{C}=u_{k}^{C}$. Hence, $\|L^{\dagger}_{k+1}r_{k}^{C}\|=\|u_{k}^{C}\|$ so
$\|u_{k}^{C}\|=\min_{x\in{\cal{K}}_{k}}\|L^{\dagger}_{k+1}(b-Ax)\|=\min_{x\in{\cal{K}}_{k}}\|L_{k+1}^{\dagger}r(x)\|$ (19)
Using (19) and the fact that $x_{k}^{G}$ is in ${\cal{K}}_{k}$ then $\|u_{k}^{C}\|\leq\|u_{k}^{G}\|$. Thus
$\|u_{k}^{C}\|\leq\|u_{k}^{G}\|=\|R_{k+1}^{-1}w_{k}^{G}\|\leq\|R_{k+1}^{-1}\|\|w_{k}^{G}\|=\|R^{-1}_{k+1}\|\|r_{k}^{G}\|,$
where the equalities in the above relation come from (18). On the other hand, and also using (18):
$\|r_{k}^{C}\|=\|L_{k+1}u_{k}^{C}\|\leq\|L_{k+1}\|\|u_{k}^{C}\|.$
Putting the above inequalities together gives the following relation:
$\|r_{k}^{C}\|$ $=$ $\|L_{k+1}u_{k}^{C}\|$
$\leq$ $\|L_{k+1}\|\|u_{k}^{C}\|$
$\leq$ $\|L_{k+1}\|\|R^{-1}_{k+1}\|\|r_{k}^{G}\|.$
Recall that $L_{k+1}$ has a QR decomposition (15) of the form $L_{k+1}=Q_{k+1}R_{k+1}$, where $Q_{k+1}$ is an orthogonal matrix. Therefore, $\|L_{k+1}\|=\|Q_{k+1}R_{k+1}\|=\|R_{k+1}\|$. This results in the following:
$\|r_{k}^{C}\|\leq\|L_{k+1}\|\|R^{-1}_{k+1}\|\|r_{k}^{G}\|=\|R_{k+1}\|\|R^{-1}_{k+1}\|\|r_{k}^{G}\|=\kappa(R_{k+1})\|r_{k}^{G}\|.$
Thus, we conclude that $\|r_{k}^{G}\|\leq\|r_{k}^{C}\|\leq\kappa(R_{k+1})\|r_{k}^{G}\|$.
3 Regularizing Properties of CMRH
As an iterative solver, CMRH produces a sequence of approximate solutions. In the early iterations, these approximations begin to converge toward the true solution, $x_{{true}}$; however, if the iterations are not stopped early, the approximate solutions eventually move away from $x_{{true}}$ as they continue to converge to the solution $x=A^{\dagger}b=A^{\dagger}(Ax_{{true}}+e)$ of the least squares problem (2), which is a poor approximation of $x_{{true}}$ due to the ill-conditioning of $A$ and the noise, $e$. This is known as semiconvergence [10]. To overcome semiconvergence, we must stop CMRH at the right time.
An analysis on the SVD of the projected matrix $H_{{k+1},k}$ from (8) provides a basis for understanding the regularizing properties of CMRH. Moreover, our hypothesis is that the approximated solutions produced by CMRH mirror regularized (filtered) solutions, so the spectral filtering properties of CMRH are also studied empirically in this section.
3.1 Singular Value Decomposition Analysis
It has been observed that the singular values of the projected matrix $H^{A}_{k+1,k}$ in (13) for GMRES (and the corresponding projected matrix for QMR [7]) tend to approximate the large singular components of $A$. This helps to explain the regularizing properties of these Krylov methods and the initial fast decay of the relative residual norms of QMR and GMRES [7].
QMR and CMRH behave similarly. Thus, we mirror the experiments in [7] by comparing the largest singular values of matrix $A$ with the singular values of the upper Hessenberg matrix $H_{{k+1},k}$ defined in (8) obtained at each iteration of CMRH. To analyze how well the singular values from $H_{{k+1},k}$ approximate those of $A$, in Figures 1 and 2 we plot the largest singular values of the full matrix $A$ using horizontal lines. The dots represent the singular values of the projected matrices at each iteration $k$ (see top plots of Figures 1 and 2). As another way of visualizing this, the bottom plots display the singular values of the upper Hessenberg matrix at different iterations of CMRH against the singular values of $A$. Each figure represents a different $1$D inverse problem; the first three (Deriv2, Heat, and Shaw) are from the Regularization Tools Package [9], and the fourth example (Spectra) is a $1$D signal restoration problem, with a matrix modeling a Gaussian blur. Specifically, the entries of $A$ are given by
$a_{ij}=\frac{1}{\varsigma\sqrt{2\pi}}\exp\left(-\frac{(i-j)^{2}}{2{\varsigma^{2}}}\right)\,,$ (20)
with $\varsigma=2$, and $x_{{true}}$ is a simulated x-ray spectrum [18].
Deriv2 Heat
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Singular values for deriv2-example 2 and heat
Shaw Spectra
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2: Singular values for shaw and spectra
From Figures 1 and 2, it seems that the quality of the singular value approximations depends on the conditioning of $A$. To investigate this, we constructed the following experiment: consider the matrix $A\in\mathbb{R}^{n\times n}$ and the vector $x_{{true}}$ generated from Spectra test problem (20). We compute the SVD of $A$ and keep the singular vectors from the orthogonal matrices $U$ and $V$ but replace the diagonal matrix containing the singular values of $A$ with a diagonal matrix $S$ whose diagonal elements are defined as:
$S^{(i)}_{kk}=e^{c_{i}*k}\quad\mbox{for}\quad k=1,...,n,$
where $c_{i}$ are entries from the vector $c=[-2,-1,-0.5,-0.25]$. For these different matrices $S^{(i)}$, we compare the largest singular values of $A^{(i)}=US^{(i)}V^{T}$ with those associated with the Hessenberg matrices obtained using the Hessenberg process associated to $A^{(i)}$ and right-hand side $b^{(i)}=A^{(i)}x_{{true}}$. The results are displayed in Figures 3 and 4. Indeed, it appears that the more ill-conditioned $A$ is, the better the singular values of the Hessenberg matrices associated to the Hessenberg process approximate the largest singular values of $A$. This result is coherent with what we observed for the previous experiments, for example in the test problems Shaw, in Figure 2, and Deriv2, in Figure 1.
Spectra at $c_{i}=-2$ Spectra at $c_{i}=-1$
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: Singular values for modified spectra
Spectra at $c_{i}=-0.5$ Spectra at $c_{i}=-0.25$
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Singular values for modified spectra
3.2 Spectral Filtering Properties
An approximated solution to (1) is said to be regularized using spectral filtering if it can be written in the form
$x=\sum_{i=1}^{n}\Phi_{i}\frac{u_{i}^{T}b}{\sigma_{i}}v_{i},$ (21)
where $u_{i}$ and $v_{i}$ correspond to left and right singular vectors of $A$, and $\sigma_{i}$ are its singular values in decreasing order of magnitude. Spectral filtering methods are a wide class of methods, including for example Tikhonov regularization, see, e.g. [10]. Moreover, note that the truncated SVD corresponds to setting $\Phi_{i}=1$ for $i$ smaller than the truncation index and $\Phi_{i}=0$ otherwise. There is also an empirical relation between Krylov methods and filtered solutions of the form (21), see, e.g. [12].
A natural question arises when studying the regularizing properties of CMRH: is there an empirical relationship between the approximate solutions computed with CMRH and a solution with spectral filtering of the form in (21)? Using (21), one can compute the empirical associated filter factors $\Phi_{i}$ to a given approximated solution at the $k$th iteration, $x_{k}$, as
$\Phi_{i}=\frac{v_{i}^{T}x_{k}}{u_{i}^{T}b}\sigma_{i},$ (22)
assuming the SVD of $A$ is available.
Consider the $1$D signal restoration problem Spectra solved using CMRH and GMRES. In Figure 5, we observe the relative error norm and its corresponding approximate solutions for three different noise levels. For each noise level, the plots in Figure 6 display the empirical filter factors for some iterations of CMRH and GMRES. These are designated with a marker in Figure 5 to illustrate the three iterations that are common for all noise levels and the last iteration before the relative error increases.
First, one can observe that for all combinations of noise levels and iterations displayed in the plots in Figure 6, the empirical filter factors for CMRH and GMRES have a very similar behavior. This is also reminiscent of the filter factors corresponding to the TSVD, which contains a sharp phase transition between filter factors being 1 and 0. Also note that this phase transition moves to the right (that is, happens at a higher value of $k$) as the iterations increase: in other words, less iterations correspond to more filtering. This is a very positive result, as it indicates that CMRH is an effective method for filtering highly oscillatory noisy components if the iterations are stopped early.
We also observe that the different noise levels (in this small regime) do not affect significantly the filter factors in the first iterations. We reiterate that in order for CMRH to mimic spectral filtering regularization, the iterative solver must terminate before the relative error increases.
Noise level 0.0001 Noise level 0.001 Noise level 0.01
Refer to caption Refer to caption Refer to caption
Figure 5: Relative error norms for the approximate solutions to test problem spectra computed using CMRH and GMRES for different noise levels. The empirical filter factors for both methods at the indicated markers can be observed in the plots in Figure 6.
iteration $k$=1 iteration $k$=4 iteration $k$=7 iteration $k$=9
Refer to caption Refer to caption Refer to caption Refer to caption
iteration $k$=1 iteration $k$=4 iteration $k$=7 iteration $k$=24
Refer to caption Refer to caption Refer to caption Refer to caption
iteration $k$=1 iteration $k$=4 iteration $k$=7 iteration $k$=34
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Empirical filter factors $\Phi_{i}$ for solutions of test problem spectra computed using GMRES and CMRH at iterations corresponding to the markers in Figure 5. The top row corresponds to left plot in Figure 5 (noise level 0.001), the middle row corresponds to the middle plot of Figure 5 (noise level 0.001), and the bottom row corresponds to the right plot in Figure 5 (noise level 0.01).
4 Hybrid Changing Minimal Residual Hessenberg Method
Consider the following optimization problem with standard Tikhonov regularization:
$\min_{x\in\mathbb{R}^{n}}\|b-Ax\|_{2}^{2}+\lambda^{2}\|x\|_{2}^{2}.$ (23)
At each iteration, we consider an approximate solution belonging to the Krylov subspace ${\cal{K}}_{k}$. Due to the lack of orthonormality of the basis vectors generated by the Hessenberg process, we solve the following least squares problem:
$\min_{x\in K_{k}}\|L^{\dagger}_{k+1}(b-Ax)\|_{2}^{2}+\lambda^{2}_{k}\|L^{\dagger}_{k}x\|_{2}^{2},$ (24)
which is an approximation of (23). Similarly to CMRH, and using the Hessenberg relationship in (8), solving (24) is equivalent to solving the following small subproblem after each iteration of the Hessenberg process:
$y_{\lambda,k}=\arg\min_{y\in\mathbb{R}^{k}}\|\beta e_{1}-H_{k+1,k}y\|_{2}^{2}+\lambda^{2}_{k}\|y\|_{2}^{2},$ (25)
where $\beta$ is the largest entry in $r_{0}$, and then project back to the original subspace using $x_{k}=x_{0}+L_{k}y_{\lambda,k}$ as seen in Algorithm 3. Note that one of the main advantages of hybrid methods is that the regularization parameter $\lambda$ in (23) can be adaptively chosen throughout the iterations: hence the subscript $\lambda_{k}$ in (24). In our case, this is due to the nature of the regularization term in (23), which implies that the projection onto the Krylov subspace is independent of the regularization parameter. Our goal is to provide an appropriate $\lambda_{k}$ at each iteration that will not cause the regularized solution $x_{k}$ to be overly smooth or oscillatory.
There exist many regularization parameter choice criteria for the Tikhonov problem, see, e.g. [11, 13, 1], and more specifically for hybrid regularization, see, e.g. [2, 8]. Moreover, since the projected problem (25) is of a small dimension, SVD based approaches of the projected matrix can be implemented to find $\lambda_{k}$ very efficiently. In this paper we analyze the results using the optimal regularization parameter with respect to the error norm and the generalized cross validation (GCV) method. These are explained in detail in Section 4.2. However, note that other regularization parameter criteria can be adapted seamlessly to this framework.
A full description of the pseudocode for H-CMRH with pivoting can be found in Algorithm 3.
Algorithm 3 Hybrid CMRH (H-CMRH)
1:Compute $r_{0}=b-Ax_{0}$
[…]
3:Determine $i_{0}$ such that $|r_{0}(i_{0})|=\|r_{0}\|_{\infty}$
4:$\beta=r_{0}(i_{0})$; $l_{1}=r_{0}/\beta$ ; $p_{1}\Leftrightarrow p_{i_{0}}$
5:for $k=1,....,m$ do
6:     $u=Al_{k}$
7:     for $j=1,....,k$ do
8:         $h_{j,k}=u(p(j))$; $u=u-h_{j,k}l_{j}$
9:     end for
10:     if $k<n$ and $u\neq 0$ then
11:         Determine $i_{0}\in\{k+1,.......,n\}$ such that $|u(p(i_{0}))|=\|u(p(k+1:n))\|_{\infty}$
12:         $h_{k+1,k}=u(p(i_{0}))$; $l_{k+1}=u/h_{k+1,k}$ ; $p_{k+1}\Leftrightarrow p_{i_{0}}$
13:     else
14:         $h_{k+1,k}=0$; Stop.
15:     end if
16:end for
17:Define the $(k+1)\times k$ Hessenberg matrix
18:Implement regularization parameter scheme ($\lambda_{k}$): optimal or GCV
19:Compute the minimizer of $\|\beta e_{1}-H_{k+1,k}y\|_{2}^{2}+\lambda^{2}_{k}\|y\|_{2}^{2}$ and $x_{k}=x_{0}+L_{k}y_{\lambda,k}$
4.1 Residual Norms for Hybrid CMRH and Hybrid GMRES
The difference between the residual norms of H-CMRH and hybrid GMRES can be bounded in a similar fashion to those of CMRH and GMRES. Since the validity of hybrid GMRES is well established, these bounds are crucial to understand the regularizing properties of H-CMRH. If we define
$\overline{L}_{k+1}=\begin{bmatrix}L_{k+1}&0\\ 0&L_{k}\end{bmatrix},$ (26)
with $L_{k},L_{k+1}$ defined by the Hessenberg relation (8), we find that if the condition number of $\overline{L}_{k+1}$ does not grow too quickly, then the residual norm associated to the solution obtained with H-CMRH is close the residual norm of the solution obtained with hybrid GMRES.
Theorem 4.1.
Let $hr_{k}^{G}$ and $hr_{k}^{C}$ be the hybrid GMRES and hybrid CMRH residuals at the kth iteration beginning with the same initial residual $r_{0}$, respectively. Then
$\|hr_{k}^{G}\|\leq\|hr_{k}^{C}\|\leq\kappa(\overline{L}_{k+1})\|hr_{k}^{G}\|,$ (27)
where $\kappa(\overline{L}_{k+1})=\|\overline{L}_{k+1}\|\|\overline{L}_{k+1}^{\dagger}\|$ is the condition number of $\overline{L}_{k+1}$.
Proof 4.2.
First, we prove the left inequality in (27). We can define the hybrid residual as a function of the solution:
$hr(x)=\begin{bmatrix}b\\ 0\end{bmatrix}-\begin{bmatrix}A\\ \lambda I\end{bmatrix}x.$
Since $x_{k}^{G}$ and $x_{k}^{C}$ are in the Krylov subspace ${\cal{K}}_{k}$, by the optimality conditions of hybrid GMRES:
$\|hr_{k}^{G}\|=\min_{x\in{\cal{K}}_{k}}\|hr(x)\|\leq\|hr(x_{k}^{C})\|=\|hr_{k}^{C}\|.$
Hence $\|hr_{k}^{G}\|\leq\|hr_{k}^{C}\|$.
Now we prove the right inequality in (27). Since $Ax$--$b\in{\cal{K}}_{k+1}$$=$range($L_{k+1}$) for any $x\in{\cal{K}}_{k}$ and $x_{k}^{G},x_{k}^{C}\in{\cal{K}}_{k}$= range($L_{k}$), we can write $hr_{k}^{C}$ and $hr_{k}^{G}$ as a linear combination of the columns of $\overline{L}_{k+1}$ defined in (26). Let $hr_{k}^{C}=\overline{L}_{k+1}u_{k}^{C}$ and $hr_{k}^{G}=\overline{L}_{k+1}u_{k}^{G}$. This implies that $u_{k}^{C}=\overline{L}_{k+1}^{\dagger}hr_{k}^{C}$ and $u_{k}^{G}=\overline{L}_{k+1}^{\dagger}hr_{k}^{G}$. Hence, $\|\overline{L}^{\dagger}_{k+1}hr_{k}^{C}\|=\|u_{k}^{C}\|$ and $\|\overline{L}^{\dagger}_{k+1}hr_{k}^{G}\|=\|u_{k}^{G}\|$. By the optimality conditions of H-CMRH:
$\|u_{k}^{C}\|=\min_{x\in{\cal{K}}_{k}}\left\|\begin{bmatrix}L^{\dagger}_{k+1}&0\\ 0&L^{\dagger}_{k}\end{bmatrix}\left(\begin{bmatrix}b\\ 0\end{bmatrix}-\begin{bmatrix}A\\ \lambda I\end{bmatrix}x\right)\right\|=\min_{x\in{\cal{K}}_{k}}\|\overline{L}^{\dagger}_{k+1}hr(x)\|.$ (28)
Using (28) and the fact that $x_{k}^{G}$ is in ${\cal{K}}_{k}$ then $\|u_{k}^{C}\|\leq\|u_{k}^{G}\|$. Thus
$\|u_{k}^{C}\|\leq\|u_{k}^{G}\|=\|\overline{L}_{k+1}^{\dagger}hr_{k}^{G}\|\leq\|\overline{L}_{k+1}^{\dagger}\|\|hr_{k}^{G}\|.$
On the other hand,
$\|hr_{k}^{C}\|=\|\overline{L}_{k+1}u_{k}^{C}\|\leq\|\overline{L}_{k+1}\|\|u_{k}^{C}\|.$
Putting the above inequalities together gives the following relation:
$\|hr_{k}^{C}\|$ $=$ $\|\overline{L}_{k+1}u_{k}^{C}\|$
$\leq$ $\|\overline{L}_{k+1}\|\|u_{k}^{C}\|$
$\leq$ $\|\overline{L}_{k+1}\|\|\overline{L}^{\dagger}_{k+1}\|\|hr_{k}^{G}\|$
$=$ $\kappa(\overline{L}_{k+1})\|hr_{k}^{G}\|\,,$
so we conclude that $\|hr_{k}^{G}\|\leq\|hr_{k}^{C}\|\leq\kappa(\overline{L}_{k+1})\|hr_{k}^{G}\|$
4.2 Regularization Parameters
To validate the potential of the CMRH method independently of the parameter choice, we use the optimal regularization parameter with respect to the error norm. This means that we minimize the following expression:
$\lambda_{k}=\arg\min_{\lambda}\|x_{\lambda,k}-x_{{true}}\|_{2}^{2},$ (29)
where $x_{\lambda,k}$ is the approximated solution at each iteration as a function of $\lambda$. Let $x_{\lambda,k}=x_{0}+L_{k}y_{\lambda,k}$. Then the minimization in (29) can be written as
$\min_{\lambda}\|x_{\lambda,k}-x_{{true}}\|_{2}^{2}=\min_{\lambda}\|(x_{0}+L_{k}y_{\lambda,k})-x_{{true}}\|_{2}^{2}.$ (30)
From (30), we use the normal equations of (25) to find a closed form expression for $y_{\lambda,k}$, such that (30) is equivalent to
$\min_{\lambda}\|(x_{0}+(L_{k}(H_{k+1,k}^{T}H_{k+1,k}+\lambda^{2}I)^{-1}H_{k+1,k}^{T}\beta e_{1}))-x_{{true}}\|_{2}^{2}$ (31)
and using the SVD of $H_{k+1,k}=U_{k}\Sigma_{k}V_{k}^{T}$ this can be further simplified to:
$\min_{\lambda}\|(x_{0}+L_{k}((V_{k}\Sigma_{k}^{T}\Sigma_{k}V_{k}^{T}+\lambda^{2}I)^{-1}V_{k}\Sigma_{k}^{T}U_{k}^{T}\beta e_{1}))-x_{{true}}\|_{2}^{2}.$ (32)
or, equivalently,
$\min_{\lambda}\|x_{\lambda,k}-x_{{true}}\|_{2}^{2}=\min_{\lambda}\|(x_{0}+L_{k}V_{k}(\Sigma_{k}^{T}\Sigma_{k}+\lambda^{2}I)^{-1}\Sigma_{k}^{T}U_{k}^{T}\beta e_{1})-x_{{true}}\|_{2}^{2}.$ (33)
This is, of course, not a reasonable regularization parameter choice criterion in practice, since it requires the knowledge of the true solution.
In a more realistic scenario, generalized cross validation (GCV) can be used to choose a suitable estimate for the regularization parameter $\lambda_{k}$ at each iteration. This scheme is a predictive statistics-based approach where prior estimates of the error norm are not needed [4]. Here, it is assumed that if a portion of arbitrary data is missing, then the regularization parameter ($\lambda_{k}$) should be able to predict the missing information. In this paper, we use the GCV function applied to the projected matrix $H_{k,k+1}$ from (8). The chosen regularization parameter using the GCV criterion minimizes the prediction error through the minimization of the GCV function:
$G(\lambda)=\frac{k\|(I-H_{k+1,k}H_{k+1,k,\lambda}^{\dagger})\beta e_{1}\|_{2}^{2}}{trace(I-H_{k+1,k}H_{k+1,k,\lambda}^{\dagger})^{2}},$ (34)
where $H_{k+1,k,\lambda}^{\dagger}=(H_{k+1,k}^{T}H_{k+1,k}+\lambda^{2}I)^{-1}H_{k+1,k}^{T}.$ Using the SVD $H_{k+1,k}=U_{k}\Sigma_{k}V_{k}^{T}$, (34) can be rewritten as:
$G(\lambda)=\displaystyle\frac{k\beta^{2}\left(\displaystyle\sum_{i=1}^{k}\left(\frac{\lambda^{2}}{\sigma_{i}^{2}+\lambda^{2}}[U^{T}_{k}e_{1}]_{i}\right)^{2}+\left(\left[U^{T}_{k}e_{1}\right]_{k+1}\right)^{2}\right)}{\left(1+\displaystyle\sum_{i=1}^{k}\frac{\lambda^{2}}{\sigma_{i}^{2}+\lambda^{2}}\right)^{2}},$ (35)
and $\lambda_{k}=\arg\min_{\lambda}G(\lambda)\,.$
4.2.1 Stopping Criterion
We must determine a suitable criterion to stop the iterations of H-CMRH. Similar to [4], we would like to find a stopping iteration $k$ that minimizes:
$\frac{n\|(I-AA_{k}^{\dagger})b\|_{2}^{2}}{(trace((I-AA_{k}^{\dagger}))^{2}},$ (36)
where $A_{k}^{\dagger}$ is defined considering the approximate solution produced by hybrid CMRH:
$x_{k}=L_{k}y_{\lambda,k}=L_{k}(H^{T}_{k+1,k}H_{k+1,k}+\lambda^{2}I)^{-1}H_{k+1,k}^{T}L_{k+1}^{\dagger}b\equiv A^{\dagger}_{k}b.$
Since $L_{k}$ does not have orthonormal columns, it is not straightforward to compute (36) exactly. However, it can be approximated using a formal parallelism with the correspondent derivations for other Krylov methods, see, e.g. [4]. This involves
$n\|(I-AA^{\dagger}_{k})b\|_{2}^{2}$$\approx$ $n\|L_{k+1}^{\dagger}(I-AA^{\dagger}_{k})b\|_{2}^{2}$
$=$ $n\|(I-H_{k+1,k}(H^{T}_{k+1,k}H_{k+1,k}+\lambda_{k}^{2}I)^{-1}H_{k+1,k}^{T})L_{k+1}^{\dagger}b\|_{2}^{2}.$
where $L_{k+1}^{\dagger}AL_{k}=L_{k+1}^{\dagger}L_{k+1}H_{k+1,k}=H_{k+1,k}$ and $L_{k+1}^{\dagger}Ib=\beta e_{1}$. Using the SVD of $H_{k+1,k}$, the previous expression can be rewritten as:
$n\|{L_{k+1}^{\dagger}}(I-AA^{\dagger}_{k})b\|_{2}^{2}=n\beta^{2}\displaystyle\left(\sum_{i=1}^{k}\left(\frac{\lambda_{k}^{2}}{\sigma_{i}^{2}+\lambda_{k}^{2}}[U_{k}^{T}e_{1}]_{i}\right)^{2}+\left([U_{k}^{T}e_{1}]_{k+1}\right)^{2}\right).$ (37)
The denominator of (36) can be similarly approximated by:
$\left(trace(I-AA_{k}^{\dagger})\right)^{2}\approx\left((n-k)+\displaystyle\sum_{i=1}^{k}\frac{\lambda_{k}^{2}}{\sigma_{i}^{2}+\lambda_{k}^{2}}\right)^{2}.$ (38)
Thus, (36) can be approximated by the following:
$\hat{G}(k)=\displaystyle\frac{n\beta^{2}\displaystyle\left(\sum_{i=1}^{k}\left(\frac{\lambda_{k}^{2}}{\sigma_{i}^{2}+\lambda_{k}^{2}}[U_{k}^{T}e_{1}]_{i}\right)^{2}+\left([U_{k}^{T}e_{1}]_{k+1}\right)^{2}\right)}{\left((n-k)+\displaystyle\sum_{i=1}^{k}\frac{\lambda_{k}^{2}}{\sigma_{i}^{2}+\lambda_{k}^{2}}\right)^{2}}\,.$ (39)
$\hat{G}(k)$ is used to determine the stopping iteration. The algorithm is terminated when the difference between values is very small:
${\left|\frac{\hat{G}(k+1)-\hat{G}(k)}{\hat{G}(1)}\right|<tol.}$ (40)
5 Numerical Results
We now demonstrate the effectiveness of H-CMRH in comparison to hybrid GMRES.
5.1 Results on various test problems
To illustrate the behavior of H-CMRH, we use three different test problems: PRblur, PRblurshake, and PRblurspeckle. PRblur generates data for the use of image deblurring. This function can generate a variety of image deblurring test problems, like PRblurshake, which contains a random motion (shaking) and PRblurspeckle, which models an atmospheric blur. These problems are 2D Deblurring Problems from the IRtools package [6] involving images with $256\times 256$ pixels, corresponding to $65536\times 65536$ matrices $A$, and noise level
$\frac{\|e\|}{\|Ax_{{true}}\|}=10^{-2}\,.$
First, we compared the blurred images with the reconstructed images that were produced by H-CMRH with the GCV scheme (see Figure 7).
PRblur PRblurshake PRblurspeckle
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Blurred and reconstructed images using H-CMRH
Next, we analyzed the relative error between H-CMRH and hybrid GMRES with the GCV and optimal schemes. The star represents where the algorithm stops based on the GCV stopping criteria in (39) and (40).
PRblur PRblurshake PRblurspeckle
Refer to caption Refer to caption Refer to caption
Figure 8: Relative Error for 2D deblurring problems using H-CMRH and hybrid gmres.
From Figures 7 and 8, we observe that over time H-CMRH and hybrid GMRES behave in almost an identical manner for PRblurshake and PRblurspeckle. For PRblur, hybrid GMRES does outperform H-CMRH. However, note that these are comparable results, and H-CMRH is an inherent inner-product free method. As proven in Section 4, the test results shown in Figures 7 and 8 illustrate that H-CMRH and hybrid GMRES have similar regularizing properties.
6 Conclusion
In this paper, we have analyzed the regularizing properties of CMRH and introduced a new hybrid regularization method, H-CMRH, for solving large scale ill-posed problems. It appears that H-CMRH is as competitive as hybrid GMRES in stabilizing semicovergence. This approach could be potentially useful for parallel computing and mixed precision problems as it does not require any inner products.